# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ 4.3 Proof of Proposition 1
$L_{\operatorname{ap}}^{U}f:=-\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle f,Q\overline{\psi}^{U}_{j}\rangle_{\Omega}\Pi(Q\overline{\psi}^{U}_{j}+r_{j}^{U}).$ (4.18)
@@
$U_{2}=G^{U}U_{1}:=\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle U_{1},Q\overline{\psi}_{j}^{U}\rangle_{\Omega}\overline{\phi}_{j}^{U},$ (4.22)
@@
$L_{1}^{U}U_{1}^{\ast}+(L_{2}^{U}-L_{\operatorname{ap}}^{U})U_{1}^{\ast}=-\Pi((U_{1}+G^{U}U_{1})\cdot\nabla(U_{1}+G^{U}U_{1}))-\Pi\mathrm{E}^{U}.$
@@
As a consequence of Lax-Milgram applied to the bilinear form (4.25), we know that there exists $U_{1}^{\ast}=T_{R}^{U}(U_{1})\in\mathcal{H}^{1}_{R}$ such that (4.23) holds.
@@
$\|U_{1}\cdot\nabla G^{U}U_{1}+G^{U}U_{1}\cdot\nabla G^{U}U_{1}\|_{L^{2}}\leqslant\|\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle U_{1},Q\overline{\psi}_{j}^{U}\rangle_{\Omega}U_{1}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}$
$+\|\sum_{j=1}^{m}\sum_{i=1}^{m}\overline{\lambda}_{j}^{-1}\langle U_{1},Q\overline{\psi}_{j}^{U}\rangle_{\Omega}\overline{\lambda}_{i}^{-1}\langle U_{1},Q\overline{\psi}_{i}^{U}\rangle_{\Omega}\overline{\phi}_{i}^{U}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}=:I_{4}+I_{5}$
@@
$|I_{4}|\leqslant\sqrt{\sum_{j=1}^{m}\langle U_{1},\overline{\lambda}_{j}^{-1}Q\overline{\psi}_{j}^{U}\rangle_{\Omega}^{2}}\sqrt{\sum_{j=1}^{m}\|\nabla\overline{\phi}_{j}^{U}\|_{L^{\infty}}^{2}}\|U_{1}\|_{L^{2}}.$
@@
$|I_{5}|\leqslant\sqrt{\sum_{j=1}^{m}\langle U_{1},\overline{\lambda}_{j}^{-1}Q\overline{\psi}_{j}^{U}\rangle_{\Omega}^{2}\sum_{i=1}^{m}\langle U_{1},\overline{\lambda}_{i}^{-1}Q\overline{\psi}_{i}^{U}\rangle_{\Omega}^{2}}\sqrt{\sum_{j=1}^{m}\sum_{i=1}^{m}\|\overline{\phi}_{i}^{U}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}^{2}}.$
@@
Notice that $B^{R}_{x_{0}^{U},x_{1}^{U}}$ is bounded in $\mathcal{H}^{1}_{R}$, and by Rellich–Kondrachov embedding theorem we know that it is a compact subset of $L^{4}$. We apply Schauder fixed-point theorem to the topology induced by $L^{4}$. Since $T^{U}_{R}$ is continuous, maps the compact subset $B^{R}_{x_{0}^{U},x_{1}^{U}}$ to itself, there exists a fixed point $U_{1}^{R}$ such that by (4.23), for any $g\in\mathcal{H}^{1}_{R}$, we have
$\langle L_{1}^{U}U_{1}^{R}+(L_{2}^{U}-L_{\operatorname{ap}}^{U})U_{1}^{R},g\rangle=\langle-(U_{1}^{R}+G^{U}U_{1}^{R})\cdot\nabla(U_{1}^{R}+G^{U}U_{1}^{R})-\mathrm{E}^{U},g\rangle,$ (4.31)
@@ 4.4 Proof of Proposition 2
We work on Proposition 2 following a similar argument to the previous subsection assuming that Proposition 1 holds, and we only highlight the key changes. We construct the finite-rank approximation operator $L_{\operatorname{ap}}^{v}$ of $L_{2}^{v}$ as:
$L_{\operatorname{ap}}^{v}f:=-\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle f,Q\overline{\psi}_{j}^{v}\rangle_{\Omega}P_{\overline{v}}\Pi(Q\overline{\psi}_{j}^{v}+r_{j}^{v}).$
@@
$\langle L_{1}^{v}v_{1}+(L_{2}^{v}-L_{\operatorname{ap}}^{v})v_{1},v_{1}\rangle$
$\geqslant$ $(\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}\|\nabla U\|_{L^{2}}-|\lambda^{\text{con}}|-M_{2}^{v})\|v_{1}\|_{L^{2}}^{2}+(1-2K_{4}\|\nabla U\|_{L^{2}}-M_{1}^{v})\|\nabla v_{1}\|_{L^{2}}^{2}.$ (4.32)
@@
$v_{2}=G^{v}v_{1}:=\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle v_{1},Q\overline{\psi}_{j}^{v}\rangle_{\Omega}\overline{\phi}_{j}^{v}.$ (4.33)
@@
for all $g\in\mathcal{H}^{1}_{R}$. We know by (4.32) that the left-hand side induces a coercive bilinear form on $\mathcal{H}^{1}_{R}$. As in the case for $U$, we know it is a bounded bilinear form. $N^{v}(v_{1}+G^{v}v_{1})-\widehat{\mathrm{E}}^{v}\in L^{2}$ trivially and by Lax-Milgram, we know that there exists $v_{1}^{\ast}:=T_{R}^{v}(v_{1})\in\mathcal{H}^{1}_{R}$ such that (4.34) holds.
@@
We have the following nonlinear estimate by (4.8), (4.33), Cauchy-Schwarz inequality, and relating the projection to the Gram matrix as in the proof of (4.27):
$|\lambda^{\text{lin}}(v_{1}+G^{v}v_{1})(v_{1}+G^{v}v_{1}),v_{1}^{\ast}\rangle|\leqslant M_{0}(\|v_{1}\|_{L^{2}}+\|G^{v}v_{1}\|_{L^{2}})^{2}\|v_{1}^{\ast}\|_{L^{2}}\leqslant M_{3}^{v}\|v_{1}\|_{L^{2}}^{2}\|v_{1}^{\ast}\|_{L^{2}},$
@@
The residual term $\widehat{\mathrm{E}}^{v}$ satisfies trivially:
$|\langle\widehat{\mathrm{E}}^{v},v_{1}^{\ast}\rangle|\leqslant(\varepsilon^{v}+\|U\|_{L^{2}}\|\nabla\overline{v}\|_{L^{\infty}}+\|\nabla U\|_{L^{2}}\|\overline{v}\|_{L^{\infty}})\|v_{1}^{\ast}\|_{L^{2}}.$
We thus collect the following $H^{1}$ estimate by (4.34) and (4.32):
$(\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}\|\nabla U\|_{L^{2}}-|\lambda^{\text{con}}|-M_{2}^{v})\|v_{1}^{\ast}\|_{L^{2}}^{2}+(1-2K_{4}\|\nabla U\|_{L^{2}}-M_{1}^{v})\|\nabla v_{1}^{\ast}\|_{L^{2}}^{2}$
$\leqslant$ $(\varepsilon^{v}+\|U\|_{L^{2}}\|\nabla\overline{v}\|_{L^{\infty}}+\|\nabla U\|_{L^{2}}\|\overline{v}\|_{L^{\infty}}+M_{3}^{v}\|v_{1}\|_{L^{2}}^{2})\|v_{1}^{\ast}\|_{L^{2}}.$
@@
$\|U\|_{L^{2}}\leqslant x_{0}^{U},\quad\|\nabla U\|_{L^{2}}\leqslant x_{1}^{U}.$
@@
$M_{4}^{v}:=(\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}x_{1}^{U}-|\lambda^{\operatorname{con}}|-M_{2}^{v})^{2}-4M_{3}^{v}(\varepsilon^{v}+x_{0}^{U}\|\nabla\overline{v}\|_{L^{\infty}}+x_{1}^{U}\|\overline{v}\|_{L^{\infty}})>0.$
@@
$x_{0}^{v}:=\frac{2(\varepsilon^{v}+x_{0}^{U}\|\nabla\overline{v}\|_{L^{\infty}}+x_{1}^{U}\|\overline{v}\|_{L^{\infty}})}{\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}x_{1}^{U}-|\lambda^{\operatorname{con}}|-M_{2}^{v}+\sqrt{M_{4}^{v}}},$
$x_{1}^{v}:=\sqrt{\frac{\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}x_{1}^{U}-|\lambda^{\operatorname{con}}|-M_{2}^{v}}{1-2K_{4}x_{1}^{U}-M_{1}^{v}}}\frac{x_{0}^{v}}{2}.$
@@
$\langle L_{1}^{v}\delta v^{\ast}+(L_{2}^{v}-L_{\operatorname{ap}}^{v})\delta v^{\ast},\delta v^{\ast}\rangle=-\langle\lambda^{\text{lin}}(\widehat{v}_{1}+G^{v}\widehat{v}_{1})(\delta v+G^{v}\delta v)+\lambda^{\text{lin}}(\delta v+G^{v}\delta v)(v_{1}+G^{v}v_{1}),\delta v^{\ast}\rangle,$ (4.35)
@@
By (4.35), (4.32), Sobolev embedding of $H^{1}$ into $L^{4}$ into $L^{2}$, and the a priori bound of $H^{1}$ norm, we conclude that
@@
$\langle L_{1}^{v}v_{1}^{R}+(L_{2}^{v}-L_{\operatorname{ap}}^{v})v_{1}^{R},g\rangle=\langle-N^{v}(v_{1}^{R}+G^{v}v_{1}^{R})-\widehat{\mathrm{E}}^{v},g\rangle,$
@@
Notice that the divergence-free condition and orthogonality condition are preserved in the weak limit. Denote $v_{1}^{\infty}\in\mathcal{H}^{1}$ as the weak limit for a subsequence of $v_{1}^{R}$. For any test function $\varphi\in C^{\infty}_{c}\cap\mathcal{H}^{1}$, we have

### v2

@@ 4.3 Proof of Proposition 1
$L_{\operatorname{ap}}^{U}f:=-\sum_{j=1}^{m}(\overline{\lambda}_{j}^{U})^{-1}\langle f,Q\overline{\psi}^{U}_{j}\rangle_{\Omega}\Pi(Q\overline{\psi}^{U}_{j}+r_{j}^{U}).$ (4.19)
@@
$U_{2}=G^{U}U_{1}:=\sum_{j=1}^{m}(\overline{\lambda}_{j}^{U})^{-1}\langle U_{1},Q\overline{\psi}_{j}^{U}\rangle_{\Omega}\overline{\phi}_{j}^{U},$ (4.23)
@@
$L_{1}^{U}U_{1}^{\ast}+(L_{2}^{U}-L_{\operatorname{ap}}^{U})U_{1}^{\ast}=-\Pi((U_{1}+G^{U}U_{1})\cdot\nabla(U_{1}+G^{U}U_{1}))-\Pi\mathrm{E}^{U},$
@@
As a consequence of the Lax-Milgram theorem applied to the bilinear form (4.26), we know that there exists $U_{1}^{\ast}=T_{R}^{U}(U_{1})\in\mathcal{H}^{1}_{R}$ such that (4.24) holds.
@@
$\|U_{1}\cdot\nabla G^{U}U_{1}+G^{U}U_{1}\cdot\nabla G^{U}U_{1}\|_{L^{2}}\leqslant\|\sum_{j=1}^{m}(\overline{\lambda}_{j}^{U})^{-1}\langle U_{1},Q\overline{\psi}_{j}^{U}\rangle_{\Omega}U_{1}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}$
$+\|\sum_{j=1}^{m}\sum_{i=1}^{m}(\overline{\lambda}_{j}^{U})^{-1}\langle U_{1},Q\overline{\psi}_{j}^{U}\rangle_{\Omega}(\overline{\lambda}_{i}^{U})^{-1}\langle U_{1},Q\overline{\psi}_{i}^{U}\rangle_{\Omega}\overline{\phi}_{i}^{U}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}=:I_{4}+I_{5}$
@@
$|I_{4}|\leqslant\sqrt{\sum_{j=1}^{m}\langle U_{1},(\overline{\lambda}_{j}^{U})^{-1}Q\overline{\psi}_{j}^{U}\rangle_{\Omega}^{2}}\sqrt{\sum_{j=1}^{m}\|\nabla\overline{\phi}_{j}^{U}\|_{L^{\infty}}^{2}}\bigg(\|U_{1}\|_{L^{2}}\bigg).$
@@
$|I_{5}|\leqslant\sqrt{\sum_{j=1}^{m}\langle U_{1},(\overline{\lambda}_{j}^{U})^{-1}Q\overline{\psi}_{j}^{U}\rangle_{\Omega}^{2}\sum_{i=1}^{m}\langle U_{1},(\overline{\lambda}_{i}^{U})^{-1}Q\overline{\psi}_{i}^{U}\rangle_{\Omega}^{2}}\sqrt{\sum_{j=1}^{m}\sum_{i=1}^{m}\|\overline{\phi}_{i}^{U}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}^{2}}.$
@@
Estimates of $U_{2}$.
In the argument above, we derived an estimate for $U_{1}$. However, this does not yet yield a bound for $U$, since $U=U_{1}+U_{2}$. For $U_{2}$, its definition implies the following bounds:
$\|U_{2}\|_{L^{2}}$ $\leqslant$ $\sum_{j=1}^{m}|(\overline{\lambda}_{j}^{U})^{-1}\langle U_{1},Q\overline{\psi}_{j}^{U}\rangle_{\Omega}|\|\overline{\phi}_{j}^{U}\|_{L^{2}}$ (4.29)
$\leqslant$ $\sqrt{\sum_{j=1}^{m}(\overline{\lambda}_{j}^{U})^{-2}|\langle U_{1},Q\overline{\psi}_{j}^{U}\rangle_{\Omega}|^{2}}\sqrt{\sum_{j=1}^{m}\|\overline{\phi}_{j}^{U}\|_{L^{2}}^{2}}$
$\leqslant$ $\|G_{Q^{2}}^{U}\|\sqrt{\sum_{j=1}^{m}\|\overline{\phi}_{j}^{U}\|_{L^{2}}^{2}}\|U_{1}\|_{L^{2}},$
and similarly,
$\|\nabla U_{2}\|_{L^{\infty}}\leqslant\|G_{Q^{2}}^{U}\|\sqrt{\sum_{j=1}^{m}\|\nabla\overline{\phi}_{j}^{U}\|_{L^{\infty}}^{2}}\|U_{1}\|_{L^{2}}.$ (4.30)
Here, for convenience, we estimate the $L^{\infty}$ norm of $\nabla U_{2}$ rather than the $L^{2}$ norm. This allows us to sharpen some of the previous estimates. For instance, (4.7) becomes
$|\langle v\cdot\nabla U,v\rangle|\leqslant|\langle v\cdot\nabla U_{1},v\rangle|+|\langle v\cdot\nabla U_{2},v\rangle|\leqslant 2K_{4}\|\nabla U_{1}\|_{L^{2}}(\|v\|_{L^{2}}^{2}+\|\nabla v\|_{L^{2}}^{2})+\|\nabla U_{2}\|_{L^{\infty}}\|v\|_{L^{2}}^{2}.$
Accordingly, the coercivity estimate for the linear operator in (4.8) becomes
$\langle L_{1}^{v}v,v\rangle\geqslant\left(\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}\|\nabla U_{1}\|_{L^{2}}-\|\nabla U_{2}\|_{L^{\infty}}-|\lambda^{\operatorname{con}}|\right)\|v\|_{L^{2}}^{2}+(1-2K_{4}\|\nabla U_{1}\|_{L^{2}})\|\nabla v\|_{L^{2}}^{2}.$ (4.31)
Similarly, the residual defined in (4.4) satisfies
$|\langle\hat{\mathrm{E}}^{v},v_{1}^{\ast}\rangle|=|\langle\Pi\mathrm{E}^{v}+\Pi(U\cdot\nabla\overline{v}+\overline{v}\cdot\nabla U)-\lambda^{\text{con}}\overline{v},v_{1}^{\ast}\rangle|\leqslant(\varepsilon^{v}+\varepsilon_{5})\|v_{1}^{\ast}\|_{L^{2}},$ (4.32)
where via the orthogonality $\langle v_{1}^{\ast},\overline{v}\rangle=0$ (4.6), we have
$\varepsilon_{5}:=(\|U_{1}\|_{L^{2}}+\|U_{2}\|_{L^{2}})\|\nabla\overline{v}\|_{L^{\infty}}+\|\nabla U_{1}\|_{L^{2}}\|\overline{v}\|_{L^{\infty}}+\|\nabla U_{2}\|_{L^{\infty}}.$ (4.33)
Since $\|\overline{v}\|_{L^{2}}=1$, the final term does not carry a factor of $\|\overline{v}\|_{L^{2}}$.
@@
Notice that $B^{R}_{x_{0}^{U},x_{1}^{U}}$ is bounded in $\mathcal{H}^{1}_{R}$, and by the Rellich–Kondrachov embedding theorem we know that it is a compact subset of $L^{4}$. We apply the Schauder fixed-point theorem to the topology induced by $L^{4}$. Since $T^{U}_{R}$ is continuous, maps the compact subset $B^{R}_{x_{0}^{U},x_{1}^{U}}$ to itself, there exists a fixed point $U_{1}^{R}$ such that by (4.24), for any $g\in\mathcal{H}^{1}_{R}$, we have
$\langle L_{1}^{U}U_{1}^{R}+(L_{2}^{U}-L_{\operatorname{ap}}^{U})U_{1}^{R},g\rangle=\langle-(U_{1}^{R}+G^{U}U_{1}^{R})\cdot\nabla(U_{1}^{R}+G^{U}U_{1}^{R})-\mathrm{E}^{U},g\rangle.$ (4.37)
@@ 4.4 Proof of Proposition 2
We work on Proposition 2 following an argument similar to the previous subsection by assuming that Proposition 1 holds, and we only highlight the key changes. We construct the finite-rank approximation operator $L_{\operatorname{ap}}^{v}$ of $L_{2}^{v}$ as:
$L_{\operatorname{ap}}^{v}f:=-\sum_{j=1}^{m}(\overline{\lambda}_{j}^{v})^{-1}\langle f,Q\overline{\psi}_{j}^{v}\rangle_{\Omega}P_{\overline{v}}\Pi(Q\overline{\psi}_{j}^{v}+r_{j}^{v}).$
@@
$\langle L_{1}^{v}v_{1}+(L_{2}^{v}-L_{\operatorname{ap}}^{v})v_{1},v_{1}\rangle\geqslant(\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}\|\nabla U_{1}\|_{L^{2}}-\|\nabla U_{2}\|_{L^{\infty}}-|\lambda^{\text{con}}|-M_{2}^{v})\|v_{1}\|_{L^{2}}^{2}$
$+$ $(1-2K_{4}\|\nabla U_{1}\|_{L^{2}}-M_{1}^{v})\|\nabla v_{1}\|_{L^{2}}^{2}.$ (4.38)
@@
$v_{2}=G^{v}v_{1}:=\sum_{j=1}^{m}(\overline{\lambda}_{j}^{v})^{-1}\langle v_{1},Q\overline{\psi}_{j}^{v}\rangle_{\Omega}\overline{\phi}_{j}^{v}.$ (4.39)
@@
for all $g\in\mathcal{H}^{1}_{R}$. We know by (4.38) that the left-hand side induces a coercive bilinear form on $\mathcal{H}^{1}_{R}$. As in the case for $U$, we know it is a bounded bilinear form. We can show $N^{v}(v_{1}+G^{v}v_{1})-\widehat{\mathrm{E}}^{v}\in L^{2}$ easily and by Lax-Milgram, we know that there exists $v_{1}^{\ast}:=T_{R}^{v}(v_{1})\in\mathcal{H}^{1}_{R}$ such that (4.40) holds.
@@
We first estimate of $\lambda^{\operatorname{lin}}$ and $\lambda^{\operatorname{con}}$, in a way similar to (4.31):
$|\lambda^{\operatorname{con}}|\leqslant\|\overline{v}\cdot\nabla U\|_{L^{2}}\leqslant\|\nabla U_{1}\|_{L^{2}}\|\overline{v}\|_{L^{\infty}}+\|\nabla U_{2}\|_{L^{\infty}}.$
For $\lambda^{\operatorname{lin}}$, we derive the following bound by integration by parts:
$|\lambda^{\operatorname{lin}}(v)|$ $=$ $|\langle\frac{1}{2}x\cdot\nabla\bar{v}-\Delta\bar{v}-\tilde{U}\cdot\nabla\bar{v}+\bar{v}\cdot\nabla^{T}\tilde{U},v\rangle|\leqslant M_{0}\|v\|_{L^{2}}$
$M_{0}$ $:=$ $\left\|P_{\bar{v}}(\frac{1}{2}x\cdot\nabla\bar{v}-\Delta\bar{v}-\bar{U}\cdot\nabla\bar{v}+\bar{v}\cdot\nabla^{T}\bar{U})\right\|_{L^{2}}+\varepsilon_{5},$ (4.41)
where we use the orthogonality of $v$ to $\bar{v}$ and the definition of $P_{\bar{v}}$ and $\varepsilon_{5}$ is defined by (4.33). By the same argument as $U_{2}$, $v_{2}$ is bounded by
$\|v_{2}\|_{L^{2}}\leqslant\|v_{1}\|_{L^{2}}\|G_{Q^{2}}^{v}\|\sqrt{\sum_{j=1}^{m}\|\overline{\phi}_{j}^{v}\|_{L^{2}}^{2}}.$
Then we have the following nonlinear estimate by (4.41), (4.39), the Cauchy-Schwarz inequality, and relating the projection to the Gram matrix as in the proof of (4.28):
$|\langle N^{v}(v_{1}+G^{v}v_{1}),v_{1}^{\ast}\rangle|=|\langle\lambda^{\text{lin}}(v_{1}+v_{2})(v_{1}+v_{2}),v_{1}^{\ast}\rangle|\leqslant M_{0}(\|v_{1}\|_{L^{2}}+\|v_{2}\|_{L^{2}})^{2}\|v_{1}^{\ast}\|_{L^{2}}\leqslant M_{3}^{v}\|v_{1}\|_{L^{2}}^{2}\|v_{1}^{\ast}\|_{L^{2}},$ (4.42)
@@
We thus collect the following $H^{1}$ estimate by (4.32), (4.38), (4.40) and (4.42):
$(\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}\|\nabla U_{1}\|_{L^{2}}-\|\nabla U_{2}\|_{L^{\infty}}-|\lambda^{\text{con}}|-M_{2}^{v})\|v_{1}\|_{L^{2}}^{2}+(1-2K_{4}\|\nabla U_{1}\|_{L^{2}}-M_{1}^{v})\|\nabla v_{1}\|_{L^{2}}^{2}$
$\leqslant(\varepsilon^{v}+\varepsilon_{5}+M_{3}^{v}\|v_{1}\|_{L^{2}}^{2})\|v_{1}^{\ast}\|_{L^{2}}.$
@@
$\|U_{1}\|_{L^{2}}\leqslant x_{0}^{U},\quad\|\nabla U_{1}\|_{L^{2}}\leqslant x_{1}^{U}.$
We also denote the bound of $U_{2}$ as
$\|U_{2}\|_{L^{2}}\leqslant y_{0}^{U},\quad\|\nabla U_{2}\|_{L^{\infty}}\leqslant y_{1}^{U}.$
@@
$M_{4}^{v}:=(M_{5})^{2}-4M_{3}^{v}(\varepsilon^{v}+\varepsilon_{5})>0.\quad M_{5}:=\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}x_{1}^{U}-y_{1}^{U}-|\lambda^{\operatorname{con}}|-M_{2}^{v}.$
@@
$x_{0}^{v}:=\frac{2(\varepsilon^{v}+\varepsilon_{5})}{M_{5}+\sqrt{M_{4}^{v}}},$
$x_{1}^{v}:=\sqrt{\frac{M_{5}}{1-2K_{4}x_{1}^{U}-M_{1}^{v}}}\frac{x_{0}^{v}}{2}.$
@@
$\langle L_{1}^{v}\delta v^{\ast}+(L_{2}^{v}-L_{\operatorname{ap}}^{v})\delta v^{\ast},\delta v^{\ast}\rangle=-\langle\lambda^{\text{lin}}(\widehat{v}_{1}+G^{v}\widehat{v}_{1})(\delta v+G^{v}\delta v)+\lambda^{\text{lin}}(\delta v+G^{v}\delta v)(v_{1}+G^{v}v_{1}),\delta v^{\ast}\rangle.$ (4.43)
@@
By (4.43), (4.38), the Sobolev embedding of $H^{1}$ into $L^{4}$ into $L^{2}$, and the a priori bound of the $H^{1}$ norm, we conclude that
@@
$\langle L_{1}^{v}v_{1}^{R}+(L_{2}^{v}-L_{\operatorname{ap}}^{v})v_{1}^{R},g\rangle=\langle-N^{v}(v_{1}^{R}+G^{v}v_{1}^{R})-\widehat{\mathrm{E}}^{v},g\rangle.$
@@
Notice that the divergence-free condition and orthogonality condition are preserved in the weak limit. Denote $v_{1}^{\infty}\in\mathcal{H}^{1}$ as the weak limit for a subsequence of $v_{1}^{R}$ as $R$ tends to infinity. For any test function $\varphi\in C^{\infty}_{c}\cap\mathcal{H}^{1}$, we have